OpenStudy (anonymous):

2/3 + 3/4 + 1/6 = ??

OpenStudy (anonymous):

Help??

OpenStudy (anonymous):

Get a common denominator of 12 for all the fractions. So, 2/3 becomes 8/12 because you multiple both the top and bottom by 4 to get 12 on bottom. 3/4 becomes 9/12, by multipling by 3. and 1/6 becomes 2/12. So, 8/12 + 9/12 + 2/12= 19/12

OpenStudy (gw2011):

You need to find a common denominator. In this case the common denominator is 12. So you get the following: 2/3=8/12 3/4=9/12 1/6=2/12 Now add these and you get: 8/12+9/12+2/12=19/12